OpenStudy (yadu123):

Consider the equation for the combustion of methanol: CH3OH + 3O2 2CO2 + 4H2O. If 265g of CO2 is produced in this reaction, what mass of methanol was required? (assume excess oxygen)

OpenStudy (yadu123):

this is chem but nobody was on that server

OpenStudy (yadu123):

and i have a test tommorow

OpenStudy (yadu123):

hello

OpenStudy (anonymous):

The coefficients of each reactant/product represents the number of moles of the substance needed in the reaction. This means you'll have to convert from mass to moles, then back to mass.

OpenStudy (anonymous):

\[\begin{array}{c|c} \text{molecule}&\text{molar mass }\left(\frac{\text{g}}{\text{mol}}\right)\\ \hline \text{CH}_3\text{OH}&12+4\times1+16=32\\ \text{O}_2&2\times16=32\\ \text{CO}_2&12+2\times16=88\\ \text{H}_2\text{O}&2\times1+16=18 \end{array}\] So, if 256 g of \(\text{CO}_2\) were produced, this gives \[256\text{ g CO}_2\times\frac{1\text{ mol CO}_2}{44\text{ g CO}_2}\times\frac{1\text{ mol CH}_3\text{OH}}{2\text{ mol CO}_2}\times\frac{32\text{ g CH}_3\text{OH}}{1\text{ mol CH}_3\text{OH}}=\cdots\text{ g CH}_3\text{OH}\]